# Discovery of three magnetic He-sdOs with SALT

### v1

@@ 5 Galactic kinematics
Galactic space velocities for the confirmed magnetic He-sdOs were calculated from their radial velocity (Table 1) and parallax and proper motion provided by Gaia DR3. Their Galactic orbits were computed with galpy [^8] http://github.com/jobovy/galpy, a python package for Galactic-dynamic calculations (Bovy, 2015). The resulting kinematic parameters are summarised in Table 4 while Fig. 12 shows the projection of the computed trajectories to the vertical distance ($z$) against Galactic radius ($R$) plane. The stars were roughly classified as Galactic thin disk, thick disk or halo stars by the method of Philip Monai et al. (2024). J1256--5753 and J1444--6744 are on almost circular orbits typical of thin disk stars while J1233--6749 is likely a thick disk star, given its larger eccentricity of $e=0.31\pm 0.01$. Similar calculations for the previously discovered magnetic He-sdOs show that, of the seven known magnetic stars, four are part of the thin disk, two are thick disk stars and one star is part of the Galactic halo, J1303+2646. This suggests that magnetic He-sdO stars can be formed in both young and old Galactic populations. This is consistent with the large range of possible delay times between the merging of double He-WD binaries and their initial formation as main sequence stars (Marsh et al., 1995; Yu & Jeffery, 2011).
@@ 6 Light curves
One might expect rotational light curve variability for strongly magnetic stars with inhomogeneous fields. Magnetic fields can cause dark or bright spots on a star’s surface, resulting in variations in brightness as the star rotates. For instance, bright magnetic spots are predicted to exist in O/B/A-type main sequence stars, caused by (sub-)surface convection zones (Cantiello & Braithwaite, 2011, 2019) or chemical spots (Shultz et al., 2018; Krticka et al., 2021), and are observed at low amplitudes in A- and late B-type stars (Balona et al., 2019). Similar variability was discovered for the helium-rich early B-type HD 144941 (Jeffery & Ramsay, 2018), a magnetic main sequence star (Przybilla et al., 2021; Shultz et al., 2021). Sub-surface convection zones are also predicted to exist in He-sdO stars (Groth et al., 1985), due to the ionisation of He ii. Momany et al. (2020) observed brightness variations in several hot subdwarfs in globular clusters at periods on the order of days and attributed this variability to rotational modulation caused by magnetic spots. Krtička et al. (2024) performed a spectroscopic analysis of three of these stars, ruling out binarity as the cause of the variations and detecting $T_{\mathrm{eff}}$ variability, though the presence of magnetic spots remains uncertain. Finally, flux variations in several cool magnetic white dwarfs are proposed to be caused by star spots (Brinkworth et al., 2013).
Considering the magnetic nature of the He-sdOs discussed here, the presence of magnetic spots on these stars seems at least plausible. We therefore searched for light curves in the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) archive up to sector 70. Unfortunately, none of the new magnetic He-sdOs have been observed with TESS so far. However, TESS data are available for three of the stars showing the 4630 Å feature. We computed frequency-amplitude spectra for these stars.
@@
J1346--4025 was observed in TESS Sectors 38 and 68 and shows an approximately sinusoidal modulation at a period of 5.4 days (Fig. 13: middle). This is commensurate with a rotational modulation and neither confirms nor excludes the presence of a magnetic field.
@@ 7 Summary and discussion
We have identified three new magnetic He-sdO stars among the sample of 592 stars observed with SALT/RSS. This brings the total number of magnetic hot subdwarfs to seven stars that form a homogeneous class.
@@ 7.2 Magnetic field
All magnetic He-sdOs have observed field strengths of 200 to 500 kG and very similar surface properties with $T_{\mathrm{eff}}$ 46 500absent46500$\approx$ 46\,500≈ 46 500 K, $\log g\approx 6$ and intermediate helium abundances at $\log n(\mathrm{He})/n(\mathrm{H})$ $\approx 0$. That is to say, the magnetic sdO stars are very localised compared with the overall hot subdwarf population. A single-star origin seems unlikely for the following reasons.
@@
    The subdwarf masses of about 0.8 ${\rm M}_{\odot}$ inferred from their locations on the helium main sequence are too high to have come from post-giant branch evolution, that is following common-envelope ejection or Roche-lobe stripping. This is because helium ignition occurs when a red-giant core reaches a mass of about 0.5 ${\rm M}_{\odot}$ and the envelope must be removed close to this point if the star is to become a hot subdwarf.
@@
    Any pre-existing core magnetic field in a hot subdwarf would likely be destroyed by shell and core convection following helium ignition, as pointed out by Cantiello et al. (2016). While less convection would result from a non-degenerate He-ignition, this cannot lead to the formation of a single hot subdwarf star, given that there is no way to remove the hydrogen envelope.
@@
    Differential rotation induced by contraction of the residual star from the giant branch to become a subdwarf could lead to a dynamo-driven field. This raises the question why this effect would not have occurred in all hot subdwarfs. There is no evidence for any detectable field in any other hot subdwarfs (Landstreet et al., 2012; Mathys et al., 2012; Randall et al., 2015).
A double-star origin seems possible. The most promising scenarios include a double He-WD merge or a merge of a hybrid HeCO-WD with a He-WD (Justham et al., 2011), as already discussed by Dorsch et al. (2022) and Pelisoli et al. (2022). Such merges are able to explain their helium-enriched surfaces and apparently single nature. This origin may be able to generate the magnetic fields but still needs to address the same question: why do not all massive subdwarfs have strong fields, in particular other He-sdO stars? A finely tuned formation scenario seems to be required.
Double-degenerate merge routes involving a fossil field in either progenitor were also considered.
@@
    A fossil-field in the accretor fails because off-centre helium ignition creates convective shells after every shell flash, extending from close to surface, after the first shell flash, through to the core, once core burning is fully established. Flash-driven convection would progressively dissipate the field throughout the interior.
@@
    A fossil-field in the donor fails because the donor material is fully mixed and/or fully convective at or near the merge. Hence a fossil field cannot survive a double white dwarf merge leading to the creation of a helium burning subdwarf.
@@
This leaves the merging of two low-mass white dwarfs as the only realistic formation scenario for magnetic He-sdOs. Recently, Pakmor et al. (2024) conducted magnetohydrodynamical (MHD) simulations of the merging of two He-WDs with masses of either 0.3 + 0.25 ${\rm M}_{\odot}$ or 0.3 + 0.3 ${\rm M}_{\odot}$. They found that strong magnetic fields can be generated during both mergers. Their simulations indicate that if helium ignites off-center, strong convection will erase these fields. However, if helium ignites in the core, the fields can persist. This core ignition scenario is likely when the merging white dwarfs have similar core densities, which corresponds to nearly equal masses. While the observed magnetic He-sdOs appear to be more massive than those in the simulations, the underlying physics might be similar. Future research should use the results from these MHD simulations to develop post-merge evolutionary models, which can predict the properties of magnetic He-sdOs once they settle to the helium main sequence.
@@
Because the helium enrichment in magnetic He-sdOs is less extreme than in the more common non-magnetic He-sdOs, new models that include a treatment of the magnetic field need to be able to retain a significant amount of surface hydrogen. Our proposal is that the magnetic subdwarfs arise from the merge of a He-WD with a H+He-WD. As they merge, the H+He-WD is destroyed and fully mixed. Most of its mass condenses onto the HeWD’s surface. Differential rotation (shear) exists at the interface between the accretor and the accreted material. This initiates a toroidal magnetic field. This toroidal field evolves by a magnetic dynamo to produce a poloidal field which is then visible at the surface.
@@
A suitable double white dwarf progenitor would have a combined mass about 0.8 ${\rm M}_{\odot}$, with component masses of about 0.4 to 0.5 ${\rm M}_{\odot}$ for the He-WD and about 0.3 to 0.4 ${\rm M}_{\odot}$ for the H+He-WD. The H+He-WD could be marginally the more massive of the pair, so long as it is has the larger radius. The hydrogen content of either component cannot be too large because the location of iHe-sdOH stars on the helium main sequence requires that there be negligible hydrogen beneath the atmosphere. Potential systems are identified theoretically (Han, 1998) and observed in both the Galactic disk and halo (Brown et al., 2020). They are believed to arise as a consequence of two common-envelope ejection episodes in an intermediate mass binary, where both episodes occur when each donor star is on the red-giant branch. The condition that some hydrogen remains on the surface of either He-WD is that the common-envelope component shrinks more quickly than the orbit decays, and before all of the hydrogen is completely removed. Hall & Jeffery (2016) discussed similar cases for a combined mass of about 0.5 ${\rm M}_{\odot}$ and found very little hydrogen could survive. An important feature of the merging double white dwarf model is that the first helium-shell flash drives a strong convection zone upwards. If this reaches the surface it reveals carbon and removes residual hydrogen (Zhang & Jeffery, 2012). Such a convective pulse would also redistribute angular momentum from the original shear layer, moving the shear layer towards the surface. The absence of carbon in the spectra of magnetic He-sdOs implies that complete mixing has not occurred and that a field-generating shear layer could still exist beneath the stellar surface. Such constraints may contribute in part to the fine tuning required to model the iHe-sdOH stars.
@@
This proposal needs to be tested more fully in a number of ways.
@@
    Demonstrate that the merge leads to sufficient differential rotation in the mixed envelope to give rise to a strong magnetic field.
@@
    Determine properties, such as mass, residual hydrogen, and differential rotation, necessary for the merged system to generate a dynamo by population synthesis or other means.
@@
    Establish the progenitor properties at the merge necessary to yield 2), such as He-WD mass, H+He-WD mass and H+He-WD core mass.
It is likely that all of these issues need to be resolved if we are to understand why there is a small sweet spot for generating strong magnetic fields in hot subdwarfs. The proposed merging of a HeWD with H+He-WD by itself can not explain why the observed magnetic He-sdOs are so similar. However, Pakmor et al. (2024) suggest that only white dwarfs with similar core densities can maintain a strong surface magnetic field in the He-sdO remnant. This finding may explain the observed clustering in the HRD because similar core densities in the merging WDs would result in similar core masses for the He-sdOs.
Unfortunately, the detailed surface metal abundances of the magnetic He-sdOs discovered here are not known because of the limited quality of the available spectra. Far-UV spectra taken with the Hubble Space Telescope could solve this issue but are presently not available.
@@ 7.3 Kinematics and light curves
We further performed Galactic orbit calculations for all seven known magnetic He-sdO stars. Given that we find four stars in the thin disk, two in the thick disk and one in the Galactic halo, magnetic He-sdOs seem to exist in all Galactic populations, both old and young. Unfortunately, no TESS light curves are available for the three new He-sdOHs. Light curve variability therefore remains undetected for all seven He-sdOHs and follow-up time series photometry should be performed.
@@ Appendix A Additional material
Table 2: Summary table of the RSS spectra taken of the three new magnetic He-sdOs. The last column lists the weighted average field strength $B_{\mathrm{mean}}$ of the two-component best-fit for each spectrum, as discussed in Sect. 3.1. $A_{2}/A_{1}$ is the surface ratio of the second field component ($B_{2}$) relative to the first ($B_{1}$).
@@
Table 3: Best stellar parameters based on the combination of spectroscopic atmospheric parameters from Table 1, the Gaia parallax, and the SED’s angular diameter. Estimated systematic uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$, and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ were added in quadrature to the statistical uncertainties for the computation of the radii $R$, luminosities $L$, and masses $M$. Stellar parameters are stated as median values with one sigma uncertainties. The last column states masses obtained by interpolation in the helium main sequence of Paczyński (1971), given the uncertainties on $T_{\mathrm{eff}}$ and $L$.
@@
999 Originally from Pelisoli et al. (2022). Originally from Dorsch et al. (2022). Too far away from the helium main sequence for a sensible mass estimate.
@@
Table 4: Galactic velocities of the new and the known magnetic He-sdOs and orbital parameters including the $z$-component of angular momentum ($L_{z}$), eccentricity (e), galactic apocentre ($R_{\rm apo}$), galactic pericentre ($R_{\rm peri}$) and maximum vertical amplitude ($z_{\rm max}$). The left-handed system for the Galactic velocity components is used here, where $U$ is positive towards the Galactic Centre, $V$ is in the direction of Galactic rotation and $W$ is positive towards the North Galactic Pole. The last column shows the population classification as described in Philip Monai et al. (2024), where TH $\equiv$ thin disk, TK $\equiv$ thick disk, and H $\equiv$ halo.

### v2

@@ 5 Galactic kinematics
Galactic space velocities for the confirmed magnetic He-sdOs were calculated from their radial velocity (Table 1) and the parallax and proper motion provided by Gaia DR3. Their Galactic orbits were computed with galpy [^8] http://github.com/jobovy/galpy, a python package for Galactic-dynamic calculations (Bovy, 2015). The resulting kinematic parameters are summarised in Table 4, while Fig. A.7 shows the projection of the computed trajectories to the vertical distance ($z$) against the Galactic radius ($R$) plane. The stars are roughly classified as Galactic thin disk, thick disk, or halo stars according to the Philip Monai et al. (2024) method. J1256--5753 and J1444--6744 are on the almost circular orbits typical of thin disk stars, while J1233--6749 is likely a thick disk star given its larger eccentricity of $e=0.31\pm 0.01$. Similar calculations for the previously discovered magnetic He-sdOs show that, of the seven known magnetic stars, four are part of the thin disk, two are thick disk stars, and one ( J1303+2646) is part of the Galactic halo. This suggests that magnetic He-sdO stars can be formed in both young and old Galactic populations. This is consistent with the large range of possible delay times between the merging of double He-WD binaries and their initial formation as main sequence stars (Marsh et al., 1995; Yu & Jeffery, 2011).
@@ 6 Light curves
One might expect rotational light curve variability for strongly magnetic stars with inhomogeneous fields. Magnetic fields can cause dark or bright spots on a star’s surface, resulting in variations in brightness as the star rotates. For instance, bright magnetic spots are predicted to exist in O/B/A-type main sequence stars, caused by (sub-)surface convection zones (Cantiello & Braithwaite, 2011, 2019) or chemical spots (Shultz et al., 2018; Krticka et al., 2021), and are observed at low amplitudes in A- and late B-type stars (Balona et al., 2019). Similar variability was discovered for the helium-rich early B-type star HD 144941 (Jeffery & Ramsay, 2018), a magnetic main sequence star (Przybilla et al., 2021; Shultz et al., 2021). Sub-surface convection zones are also predicted to exist in He-sdO stars (Groth et al., 1985) due to the ionisation of He ii. Momany et al. (2020) observed brightness variations in several hot subdwarfs in globular clusters at periods of the order of days and attributed this variability to rotational modulation caused by magnetic spots. Krtička et al. (2024) performed a spectroscopic analysis of three of these stars, ruling out binarity as the cause of the variations and detecting $T_{\mathrm{eff}}$ variability, though the presence of magnetic spots remains uncertain. Finally, flux variations in several cool magnetic WDs are proposed to be caused by star spots (Brinkworth et al., 2013).
Considering the magnetic nature of the He-sdOs discussed here, the presence of magnetic spots on these stars seems at least plausible. We therefore searched for light curves in the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2014) archive up to sector 70. Unfortunately, none of the new magnetic He-sdOs have been observed with TESS so far. However, TESS data are available for three of the stars that show the 4630 Å feature. We computed frequency-amplitude spectra for these stars.
@@
J1346--4025 was observed in TESS sectors 38 and 68 and shows an approximately sinusoidal modulation with a period of 5.4 days (Fig. 13: middle). This is commensurate with a rotational modulation and neither confirms nor excludes the presence of a magnetic field.
@@ 7 Summary and discussion
We have identified three new magnetic He-sdO stars among the sample of 592 stars observed with SALT/RSS. This brings the total number of magnetic hot subdwarfs to seven, which form a homogeneous class.
@@ 7.2 Magnetic field
All magnetic He-sdOs have observed field strengths of 200 to 500 kG and very similar surface properties, with $T_{\mathrm{eff}}$ 46 500absent46500$\approx$ 46\,500≈ 46 500 K, $\log g\approx 6,$ and intermediate helium abundances at $\log n(\mathrm{He})/n(\mathrm{H})$ $\approx 0$. That is to say, the magnetic sdO stars are very localised compared with the overall hot subdwarf population. A single-star origin seems unlikely for the following reasons.
@@
    The subdwarf masses of about 0.8 ${\rm M}_{\odot}$ inferred from their locations on the helium main sequence are too high to have come from post giant branch evolution, that is, following common-envelope ejection or Roche-lobe stripping. This is because helium ignition occurs when a red-giant core reaches a mass of about 0.5 ${\rm M}_{\odot}$, and the envelope must be removed close to this point if the star is to become a hot subdwarf.
@@
    Any pre-existing core magnetic field in a hot subdwarf would likely be destroyed by shell and core convection following helium ignition, as pointed out by Cantiello et al. (2016). While less convection would result from a non-degenerate He ignition, this cannot lead to the formation of a single hot subdwarf star, given that there is no way to remove the hydrogen envelope.
@@
    Differential rotation induced by the contraction of the residual star from the giant branch to become a subdwarf could lead to a dynamo-driven field. This raises the question of why this effect would not have occurred in all hot subdwarfs. There is no evidence for any detectable field in any other hot subdwarfs (Landstreet et al., 2012; Mathys et al., 2012; Randall et al., 2015).
A double-star origin seems possible. The most promising scenarios include a double He-WD merger or a merger of a hybrid HeCO-WD with a He-WD (Justham et al., 2011), as already discussed by Dorsch et al. (2022) and Pelisoli et al. (2022). Such mergers would explain their helium-enriched surfaces and apparently single nature. This origin may be able to generate the magnetic fields but we are still left with the same question: why do not all massive subdwarfs have strong fields, in particular other He-sdO stars? A finely tuned formation scenario seems to be required.
Double-degenerate merger routes involving a fossil field in either progenitor were also considered.
@@
    A fossil field in the accretor fails because off-centre helium ignition creates convective shells after every shell flash, extending from close to surface, after the first shell flash, through to the core, once core-burning is fully established. Flash-driven convection would progressively dissipate the field throughout the interior.
@@
    A fossil field in the donor fails because the donor material is fully mixed and/or fully convective at or near the merger. Hence, a fossil field cannot survive a double WD merger leading to the creation of a helium-burning subdwarf.
@@
This leaves the merging of two low-mass WDs as the only realistic formation scenario for magnetic He-sdOs. Recently, Pakmor et al. (2024) conducted magnetohydrodynamic simulations of the merging of two He-WDs with masses of either 0.3 + 0.25 ${\rm M}_{\odot}$ or 0.3 + 0.3 ${\rm M}_{\odot}$. They find that strong magnetic fields can be generated during both mergers. Their simulations indicate that if helium ignites off-centre, strong convection will erase these fields. However, if helium ignites in the core, the fields can persist. This core ignition scenario is likely when the merging WDs have similar core densities, which corresponds to nearly equal masses. While the observed magnetic He-sdOs appear to be more massive than those in the simulations, the underlying physics might be similar. Future research should use the results from these magnetohydrodynamic simulations to develop post-merger evolutionary models, which can predict the properties of magnetic He-sdOs once they settle onto the helium main sequence.
@@
Because the helium enrichment in magnetic He-sdOs is less extreme than in the more common non-magnetic He-sdOs, new models that include a treatment of the magnetic field need to be able to retain a significant amount of surface hydrogen. Our proposal is that the magnetic subdwarfs arise from the merging of a He-WD with a H+He-WD. As they merge, the H+He-WD is destroyed and fully mixed. Most of its mass condenses onto the He-WD’s surface. Differential rotation (shear) exists at the interface between the accretor and the accreted material. This initiates a toroidal magnetic field. This toroidal field evolves via a magnetic dynamo to produce a poloidal field, which is then visible at the surface.
@@
A suitable double-WD progenitor would have a combined mass about 0.8 ${\rm M}_{\odot}$, with component masses of about 0.4 to 0.5 ${\rm M}_{\odot}$ for the He-WD and about 0.3 to 0.4 ${\rm M}_{\odot}$ for the H+He-WD. The H+He-WD could be marginally the more massive of the pair, so long as it has the larger radius. The hydrogen content of either component cannot be too large because the location of iHe-sdOH stars on the helium main sequence requires that there be negligible hydrogen beneath the atmosphere. Potential systems have been identified theoretically (Han, 1998) and observed in both the Galactic disk and halo (Brown et al., 2020). They are believed to arise as a consequence of two common-envelope ejection episodes in an intermediate-mass binary, where both episodes occur when each donor star is on the red-giant branch. The condition that some hydrogen remains on the surface of either He-WD is that the common-envelope component shrinks more quickly than the orbit decays, and before all of the hydrogen is completely removed. Hall & Jeffery (2016) discuss similar cases for a combined mass of about 0.5 ${\rm M}_{\odot}$ and find that very little hydrogen could survive. An important feature of the merging double WD model is that the first helium-shell flash drives a strong convection zone upwards. If this zone reaches the surface, it reveals carbon and removes residual hydrogen (Zhang & Jeffery, 2012). Such a convective pulse would also redistribute angular momentum from the original shear layer, moving the shear layer towards the surface. The absence of carbon in the spectra of magnetic He-sdOs implies that complete mixing has not occurred and that a field-generating shear layer could still exist beneath the stellar surface. Such constraints may contribute in part to the fine-tuning required to model the iHe-sdOH stars.
@@
This proposal needs to be tested more fully in a number of ways. We need to:
@@
    Demonstrate that the merger leads to sufficient differential rotation in the mixed envelope to give rise to a strong magnetic field.
@@
    Determine properties, such as the mass, residual hydrogen, and differential rotation, necessary for the merged system to generate a dynamo via population synthesis or other means.
@@
    Establish the progenitor properties at the merge necessary to yield (2), such as the He-WD mass, H+He-WD mass, and H+He-WD core mass.
All of these issues likely need to be resolved if we are to understand why there is a small sweet spot for generating strong magnetic fields in hot subdwarfs. The proposed merging of a He-WD with a H+He-WD by itself cannot explain why the observed magnetic He-sdOs are so similar. However, Pakmor et al. (2024) suggest that only WDs with similar core densities can maintain a strong surface magnetic field in the He-sdO remnant. This finding may explain the observed clustering in the Hertzsprung-Russell diagram because similar core densities in the merging WDs would result in similar core masses for the He-sdOs.
Unfortunately, the detailed surface metal abundances of the magnetic He-sdOs discovered here are not known, because of the limited quality of the available spectra. Far-UV spectra taken with the Hubble Space Telescope could solve this issue but are presently not available.
@@ 7.3 Kinematics and light curves
We further performed Galactic orbit calculations for all seven known magnetic He-sdO stars. Given that we find four stars in the thin disk, two in the thick disk, and one in the Galactic halo, magnetic He-sdOs seem to exist in all Galactic populations, both old and young. Unfortunately, no TESS light curves are available for the three new He-sdOHs. Light curve variability therefore remains undetected for all seven He-sdOHs, and follow-up time series photometry should be performed.
@@ Appendix A Additional material
At the editor’s request, Figs. A.1 to A.8 are available on Zenodo only.
Table 2: RSS spectra of the three new magnetic He-sdOs.
@@
999 The last column lists the weighted average field strength $B_{\mathrm{mean}}$ of the two-component best-fit for each spectrum, as discussed in Sect. 3.1. $A_{2}/A_{1}$ is the surface ratio of the second field component ($B_{2}$) relative to the first ($B_{1}$).
Table 3: Best stellar parameters based on the combination of spectroscopic atmospheric parameters from Table 1, the Gaia parallax, and the SED’s angular diameter.
@@
101010 Originally from Pelisoli et al. (2022). Originally from Dorsch et al. (2022). Too far away from the helium main sequence for a sensible mass estimate. Estimated systematic uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$, and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ were added in quadrature to the statistical uncertainties for the computation of the radii $R$, luminosities $L$, and masses $M$. Stellar parameters are stated as median values with one sigma uncertainties. The last column states masses obtained by interpolation in the helium main sequence of Paczyński (1971), given the uncertainties on $T_{\mathrm{eff}}$ and $L$.
@@
Table 4: Galactic velocities of the new and the known magnetic He-sdOs and orbital parameters.
@@
111111 The table lists the $z$-component of angular momentum ($L_{z}$), eccentricity (e), galactic apocentre ($R_{\rm apo}$), galactic pericentre ($R_{\rm peri}$) and maximum vertical amplitude ($z_{\rm max}$). The left-handed system for the Galactic velocity components is used here, where $U$ is positive towards the Galactic Centre, $V$ is in the direction of Galactic rotation and $W$ is positive towards the North Galactic Pole. The last column shows the population classification as described in Philip Monai et al. (2024), where TH $\equiv$ thin disk, TK $\equiv$ thick disk, and H $\equiv$ halo.